# Few-Shot Adversarial Prompt Learning on Vision-Language Models

### v1

@@ 4.2 Main Results
(a) Average on 11 Datasets
(b) ImageNet
(c) Caltech101
(d) Flowers102
(e) DTD
(f) Food101
(g) EuroSAT
(h) OxfordPets
(i) FGVCAircraft
(j) StanfordCars
(k) SUN397
(l) UCF101
Adversarial cross-dataset evaluation. For the cross-dataset evaluation, models are adapted on the ImageNet dataset using 16 shots and then assessed for their zero-shot adversarial robustness across 10 distinct datasets, without further downstream tuning. As shown in Table 2, our method outperforms its counterparts in 8/11 datasets and baseline in all 11 datasets. Moreover, it reveals that robust adaptation takes the cost of natural accuracy, as models obtained using various robust adaptation methods exhibit a decline in zero-shot natural accuracy performance on downstream datasets, compared to the original CLIP model.
Adversarial base-to-new generalization. We present a more challenging adversarial base-to-new generalization setting, where datasets are bifurcated into base and new subclasses. Here, models are trained with a 16-shot dataset from the base classes and are subsequently evaluated on both base and new classes. In this setting, as the number of categories in datasets is generally much smaller than the number of examples per class, models need to learn both intrinsic features within each dataset and robust representations from limited examples to achieve effective generalization on large amounts of test data.
@@
| Natural Acc (%) | ImageNet | Caltech101 | DTD | EuroSAT | OxfordPets | FGVCAircraft | Food101 | Flowers102 | StanfordCars | SUN397 | UCF101 | Average |
| PGD-100 Acc (%) | ImageNet | Caltech101 | DTD | EuroSAT | OxfordPets | FGVCAircraft | Food101 | Flowers102 | StanfordCars | SUN397 | UCF101 | Average |
4.3 More Analysis
Matching benchmark result adapted with ImageNet-1K. Beyond comparison with the baseline AdvVP under the few-shot setting, we also present its benchmark result which is adapted on the entire ImageNet with a larger prompt size (token prompt of size 200 and pad prompter of size 40) (Mao et al., 2022). In Table 3, our method matches the benchmark result with 1.25% examples from ImageNet, thus speeding up the training process by a large margin (over 97%). Furthermore, by increasing training examples from 16-shot to 32-shot and prompt depth from 9 to 12, the proposed method surpasses the previous state-of-the-art adversarial prompt tuning.
@@ 4.3 More Analysis
Trade-off between natural and adversarial robustness. Aligning with the decoupled form of classical adversarial training (Zhang et al., 2019), our prompt objective incorporates two terms that respectively ensure the generalization of natural examples and the consistency of robust representations. This motivates us to investigate the trade-off between natural and adversarial robustness, and to dynamically adjust this trade-off depending on the desired level of adversarial robustness.
@@
Natural generalization gap hinders robust adapting. We identify a failure risk in few-shot adversarial prompt learning using TeCoA loss (Mao et al., 2022), where insufficient natural generalization on the sampled dataset impedes robustness learning. Figure 3 (left) displays the loss variation during training under this setup. Under the same experimental setup using the TeCoA loss, different trials exhibit completely different trends: the curve for the failure case shows that the loss quickly ceases to decline and becomes stable shortly after training begins, whereas the loss in the normal case continues to decrease as the training progresses.
We presume that this failure stems from a lack of natural generalization ability. To confirm this, we first conduct natural tuning on the problematic few-shot dataset and then apply adversarial prompt learning. This restores the model’s robust fine-tuning performance, as evident in Figure 3 (right), where natural and robust accuracies improve significantly after natural example adaptation. Besides, we validate the learning process on the same few-shot dataset with a dual-form loss in the training objective that considers both natural and adversarial terms (red lines in Figure 3 (left)). It is revealed that this two-term loss effectively acts as a surrogate for the aforementioned two-stage method, avoiding potential failures caused by the natural generalization barrier in end-to-end training.
@@ D Additional Experimental Results
- D.1 Instability Analysis for Deep Prompt Interaction
- D.2 Case Study: Improving AdvVLP with Our Learning Objective
- D.3 Ablation for Loss Function Design
- D.4 Results on Different CLIP Architectures
- D.5 Zero-shot Adversarial Robustness under Different Perturbation Bounds
- D.6 Zero-shot Adversarial Evaluation under Auto-Attack
- D.7 Detailed Results for Adversarial Few-shot Learning
- D.8 Detailed Results for Adversarial Base-to-New Generalization
- D.9 Comparison between Adversarial Text and Vision Prompt
- E Reproducibility
- F Limitations
@@ Appendix B Related Work
Adversarial robustness. Adversarial attacks fool models by overlaying carefully designed imperceptible perturbations on input data (Szegedy et al., 2013; Goodfellow et al., 2014). In response to the susceptibility of models to such attacks, adversarial training (Goodfellow et al., 2014; Madry et al., 2017; Zhang et al., 2019, 2020; Wang et al., 2020; Wang & Wang, 2022) has emerged as one of the most effective empirical defense methods to enhance model robustness. It incorporates adversarial data into the training process and ensures the model’s predictive distribution for adversarial images closely aligns with the ground truth label. Moreover, recent advancements have seen the incorporation of contrastive learning into adversarial training (Oord et al., 2018; Chen et al., 2020a, b), which enables models to learn robust feature representations through instance discrimination tasks. As a result, models can align predictions for natural and adversarial image pairs in a self-supervised manner (Kim et al., 2020; Jiang et al., 2020; Fan et al., 2021; Yu et al., 2022c; Zhang et al., 2022a; Xu et al., 2023a, b). Additionally, there’s a growing interest in aligning predictions for adversarial image-text pairs in a text-supervised context (Mao et al., 2022; Li et al., 2023b), offering new avenues for zero-shot adversarial evaluation. Nevertheless, current research utilizes CLIP text encoding to produce static text supervision, which, although effective for clean images, may not adequately cater to the nuances of adversarial examples.
@@
Adversarial few/zero-shot classification. Adversarial training possesses a significantly larger sample complexity of robust generalization (Yu et al., 2022a), making it challenging to learn robust representations from sparse data. Existing works in adversarial few-shot classification fall into two categories: meta-learning based (Yin et al., 2018; Goldblum et al., 2020; Wang et al., 2021), which optimize an adversarial meta-learner using both clean and adversarial examples, and non-meta-learning based (Dong et al., 2022; Subramanya & Pirsiavash, 2022), employing strategies like auxiliary corrective classifiers (Dong et al., 2022; Subramanya & Pirsiavash, 2022) or reweighted mechanisms (Dong et al., 2022) for learning robust embeddings. Additionally, Yucel et al. (2020) initiated the investigation of adversarial robustness in a zero-shot learning setting, where no downstream statistics are available during training. Inspired by the successes of Vision Language Models (VLMs), recent studies (Mao et al., 2022; Zhang et al., 2023b) have unanimously chosen to incorporate semantic information from text supervision to bridge the generalization gap.
@@
Vision-language models (VLMs). Foundational VLMs (Radford et al., 2021; Jia et al., 2021; Kim et al., 2021; Yao et al., 2021; Yuan et al., 2021; Li et al., 2022a; Yu et al., 2022b; Li et al., 2023a) integrate interactions derived from image and text encodings for multi-modal pre-training. Depending on their specific objectives, VLMs can be trained through image-text contrastive learning (Radford et al., 2021; Jia et al., 2021; Yao et al., 2021; Yuan et al., 2021; Li et al., 2022a; Zhai et al., 2022; Li et al., 2023a), image-text matching (Li et al., 2022a, 2023a), and text generation (Li et al., 2022a; Yu et al., 2022b; Li et al., 2023a). Utilizing large-scale image-text datasets (e.g., 400M pairs for CLIP (Radford et al., 2021), 1B for ALIGN (Jia et al., 2021)) and end-to-end pre-training, these models acquire semantic relations between text and image features, thus exhibiting a profound understanding of open-vocabulary concepts. Consequently, VLMs have emerged as state-of-the-art solutions for various visual and vision-language tasks (Gu et al., 2021; Xu et al., 2022; Li et al., 2022b; Luo et al., 2022; Vinker et al., 2022; Chen et al., 2023). Nevertheless, some recent researches (Devillers et al., 2021; Zhang et al., 2022b) reveal that VLMs are also highly susceptible to adversarial perturbations.
@@
Prompt learning for VLMs. Prompt learning, initially introduced in the NLP community (Lester et al., 2021; Liu et al., 2023b, a), involves adapting pre-trained models by adding a small number of new learnable parameters in the input data for downstream tasks, without altering the pre-trained weights. This method stands out among other lightweight adaptation approaches due to its exceptional adaptability and flexibility. It has garnered increasing attention for adapting vision (Jia et al., 2022; Bahng et al., 2022; Bar et al., 2022) and vision-language models (Zhou et al., 2022b, a; Lu et al., 2022; Shu et al., 2022; Khattak et al., 2023a, b; Zhu et al., 2023). Specifically, in VLMs, CoOp (Zhou et al., 2022b) pioneers prompt engineering for adapting CLIP models by modeling learnable context vectors to replace hand-crafted text prompts. CoCoOp (Zhou et al., 2022a) further enhances the generalization ability of CoOp by introducing conditional prompts specific to each visual input instance. MaPLe (Khattak et al., 2023a) integrates vision and language prompts with inner synergy for cross-modality prompt learning. Two recent works, ProGrad (Zhu et al., 2023) and PromptSRC (Khattak et al., 2023b), concurrently advance the generalization of prompt learning by employing regulating constraints from zero-shot CLIP predictions to prevent the forgetting of general knowledge.
@@ Appendix C Additional Implementation Details
All experiments are conducted in an environment running PyTorch 1.10.1 and CUDA 11.3 on Python 3.8. Experiments of adversarial prompt tuning on the ImageNet dataset are carried out on a single NVIDIA RTX A40 GPU, while experiments on the other 10 datasets are performed on a single NVIDIA RTX 4090 GPU.
@@ C.1 Additional Implementation Details for Baselines
Adversarial visual prompt. We implement the adversarial visual prompt following all architectural and parameter settings in (Mao et al., 2022) for a fair comparison. In detail, we follow their code implementation to use a token-level prompt with size 5 and an image padding prompt for 30 pixels around the image. An SGD optimizer and a consine learning rate scheduler are used to train 10 epochs with an initial learning rate of 40.
Adversarial text prompt. We adopt a CoOp architecture (Zhou et al., 2022b) as our text prompt baseline and adapt learnable context vectors with adversarial examples. We typically follow (Zhou et al., 2022b) to use 16 context tokens with an additional class token appended at the end of the context vector. An SGD optimizer and a consine learning rate scheduler are used to train 200 epochs with an initial learning rate of 0.002, which aligns with the training settings in CoOp.
Adversarial multi-modal prompt. Adversarial multi-modal prompt in this work follows all the design choices as MaPLe(Khattak et al., 2023a), but are adapted with adversarial text-image contrastive loss. To sum up, it contains a token-level learnable token with size 2 in both text and visual branches in the first 9 transformer layers, and the deep prompts are coupled through a text-to-image projection. The above prompt tokens as well as the deep projections are optimized for 10 epochs with SGD optimizer and cosine learning rate scheduler from an initial learning rate of 0.0035.
@@ Appendix D Additional Experimental Results
D.1 Instability Analysis for Deep Prompt Interaction
We report an instability of generalization performance caused by the improper deep prompt interaction settings. That is, the standard cross-modal prompt interaction design, from text prompt token to image prompt token, is not plug-and-play under the setting of adversarial robustness. When natural and adversarial terms are present in a certain moderate ratio in the learning objective, the performance of the model may experience a significant decline. From Figure 4, we find that the instability intensity caused by the text-to-image design varies across different datasets, and the values of leading to this instability are also different. For instance, on some generic datasets, the performance degradation it usually brings is not significant (Figure 3(c)). However, on some fine-grained datasets, the significant performance degradation caused by this instability is unacceptable (Figure 3(b)).
To understand this, we plot the loss curve during the training process under both stable and unstable settings. As revealed in Figure 5, in unstable cases, we observe that the robust loss drops to zero early in training and remains nearly unchanged at this low level during the mid-phase, while the overall loss does not decrease as expected. This phenomenon seems to stem from the text prompt falling into a local trivial solution during optimization. Specifically, the prompt adjusts to equate natural and adversarial text-image logits, effectively nullifying the adversarial term but neglecting natural generalization, resulting in a persistently high natural loss. This training issue typically arises when the natural and robust terms in the training objective maintain a moderate ratio.
We make the following explanation for this issue: A smaller , favoring the natural term, means that reducing the robust loss rapidly in early training does not significantly lower the total loss. Conversely, When is large, due to the higher proportion of the robust term, the robust loss decreases rapidly to zero with a substantial gradient, quickly reducing the total loss. However, the robust loss does not maintain at zero because the momentum brought about by the larger gradient drives the total loss away from the local optimum.
We make a minimal refinement to prevent this instability. If we switch the deep prompt interaction to an image-to-text scenario, the text prompt can no longer be freely adapted but is instead derived from the projection of the image prompt. As the adversarial loss gradually decreases, it does not reach zero, thus avoiding the aforementioned issue.
D.2 Case Study: Improving AdvVLP with Our Learning Objective

### v2

@@ 4.2 Main Results
Adversarial base-to-new generalization. We present a more challenging adversarial base-to-new generalization setting, where datasets are bifurcated into base and new subclasses. Here, models are trained with a 16-shot dataset from the base classes and are subsequently evaluated on both base and new classes. In this setting, as the number of categories in datasets is generally much smaller than the number of examples per class, models need to learn intrinsic features within each dataset and robust representations from limited examples to effectively generalize large amounts of test data.
@@
Matching Benchmark Zero-Shot Results Adapted with ImageNet-1K. In addition to comparing with the baseline AdvVP under few-shot settings, we also benchmark against zero-shot results, where robustness is evaluated through cross-dataset evaluations. Initially adapted on ImageNet-1K, our method does not require adaptation across the entire dataset nor extensive prompt designs like the AdvVP [11], which uses embedding-level token prompts of size 200 and pixel-level pad prompters of size 40. As shown in Table 2, our method aligns with benchmark performance using just 1.25% of ImageNet-1K examples, significantly accelerating the training process by over 97%. Moreover, enhancements from 16-shot to 32-shot training and deepening prompt layers from 9 to 12 allow our method to exceed previous adversarial prompt tuning results.
@@
4.3 More Analysis
Trade-off between natural and adversarial robustness. Aligning with the decoupled form of classical adversarial training [31], our prompt objective incorporates two terms that ensure the generalization of natural examples and the consistency of robust representations. This motivates us to investigate the trade-off between natural and adversarial robustness, and to dynamically adjust this trade-off depending on the desired level of adversarial robustness.
@@ 4.3 More Analysis
Ablation for training objective design. In Section 3.4, we present our proposed novel training objective tailored for adversarial prompt learning. Our loss follows a two-term design, comprising a natural term and an adversarial term. The adversarial term further considers both the consistency and diversity of natural and adversarial features. In practice, we use KL divergence to constrain cross-modal consistency and encourage uni-modal diversity with cosine similarity. In Table 5, we present other possible designs for the loss function and conduct an ablation study under the adversarial base-to-new setting. Our method provides the best robustness across all these loss function settings.
| Natural term | Adversarial term | Base Nat Acc | Base Adv Acc | New Nat Acc | New Adv Acc | |
| Consistency | Diversity | |||||
| ✘ | TeCoA | ✘ | ||||
| ✔ | TeCoA | ✘ | ||||
| ✔ | JS | ✘ | ||||
| ✔ | KL | ✘ | ||||
| ✔ | KL | MSE | ||||
| ✔ | KL | MAE | ||||
| ✔ | KL | Cos | ||||
Instability analysis for deep prompt interaction. We report an instability of generalization performance caused by the improper deep prompt interaction, revealing that the standard cross-modal prompt interaction design, from text to image prompt token, is not plug-and-play under the setting of adversarial robustness. When natural and adversarial terms are present in a certain moderate ratio in the learning objective, the performance of the model may experience a significant decline. From Figure 4, we find that the instability intensity caused by the text-to-image design varies across different datasets, and the values of leading to this instability are also different. For instance, on some generic datasets, the performance degradation it usually brings is not significant (Figure 4(c)). However, on some fine-grained datasets, the significant performance degradation caused by this instability is unacceptable (Figure 4(b)).
To understand this, we plot the loss curve during the training process under both stable and unstable settings. As revealed in Figure 5, in unstable cases, we observe that the robust loss drops to zero early in training and remains nearly unchanged at this low level during the mid-phase, while the overall loss does not decrease as expected. This suggests the text prompt falls into a trivial local solution during optimization, equating natural and adversarial logits. This nullifies the adversarial term but overlooks natural generalization, causing consistently high natural loss. This issue typically occurs when the natural and robust terms are balanced in a moderate ratio in the training objective.
We propose a minimal refinement to prevent instability: switching the deep prompt interaction to an image-to-text scenario. Here, the text prompt is derived from the image prompt projection, limiting its adaptability. This prevents the adversarial loss from reaching zero, thus avoiding the issue.
@@ D Additional Experimental Results
- D.1 Detailed Results on Adversarial Cross-Dataset Evaluation
- D.2 Natural Generalization Gap Hinders Robust Adapting
- D.3 Incremental Changes from AdvMaPLe
- D.4 Case Study: Improving AdvVLP with Our Learning Objective
- D.5 Results on Different CLIP Architectures
- D.6 Zero-shot Adversarial Robustness under Different Perturbation Bounds
- D.7 Zero-shot Adversarial Evaluation under Auto-Attack
- D.8 Discussions on Training-time Attack Generation
- D.9 Detailed Results for Adversarial Few-shot Learning
- D.10 Detailed Results for Adversarial Base-to-New Generalization
- D.11 Comparison between Adversarial Text and Vision Prompt
- E Impact Statement
- F Reproducibility
- G Limitations
@@ Appendix B Related Work
Adversarial robustness. Adversarial attacks fool models by overlaying carefully designed imperceptible perturbations on input data [1, 2, 43]. In response to the susceptibility of models to such attacks, adversarial training [2, 44, 31, 45, 46, 47, 48, 49] has emerged as one of the most effective empirical defense methods to enhance model robustness. It incorporates adversarial data into the training process and ensures the model’s predictive distribution for adversarial images closely aligns with the ground truth label. Moreover, recent advancements have seen the incorporation of contrastive learning into adversarial training [50, 51, 52], which enables models to learn robust feature representations through instance discrimination tasks. As a result, models can align predictions for natural and adversarial image pairs in a self-supervised manner [53, 54, 55, 56, 57, 58, 59]. Additionally, there’s a growing interest in aligning predictions for adversarial image-text pairs in a text-supervised context [11, 60], offering new avenues for zero-shot adversarial evaluation. Nevertheless, current research utilizes CLIP text encoding to produce static text supervision, which, although effective for clean images, may not adequately cater to the nuances of adversarial examples.
@@
Adversarial few/zero-shot classification. Adversarial training possesses a significantly larger sample complexity of robust generalization [61], making it challenging to learn robust representations from sparse data. Existing works in adversarial few-shot classification fall into two categories: meta-learning based [62, 63, 21], which optimize an adversarial meta-learner using both clean and adversarial examples, and non-meta-learning based [22, 64], employing strategies like auxiliary corrective classifiers [22, 64] or reweighted mechanisms [22] for learning robust embeddings. Additionally, Yucel et al. [65] initiated the investigation of adversarial robustness in a zero-shot learning setting, where no downstream statistics are available during training. Inspired by the successes of Vision Language Models (VLMs), recent studies [11, 66] have unanimously chosen to incorporate semantic information from text supervision to bridge the generalization gap.
@@
Vision-language models (VLMs). Foundational VLMs [12, 13, 14, 15, 16, 17, 18, 19, 67, 68] integrate interactions derived from image and text encodings for multi-modal pre-training. Depending on their specific objectives, VLMs can be trained through image-text contrastive learning [12, 13, 15, 16, 17, 69, 19], image-text matching [17, 19], and text generation [17, 18, 19]. Utilizing large-scale image-text datasets (e.g., 400M pairs for CLIP [12], 1B for ALIGN [13]) and end-to-end pre-training, these models acquire semantic relations between text and image features, thus exhibiting a profound understanding of open-vocabulary concepts. Consequently, VLMs have emerged as state-of-the-art solutions for various visual and vision-language tasks [70, 71, 72, 73, 74, 75, 76]. Nevertheless, some recent researches [77, 78] reveal that VLMs are also highly susceptible to adversarial perturbations.
@@
Prompt learning for VLMs. Prompt learning, initially introduced in the NLP community [23, 24, 25], involves adapting pre-trained models by adding a small number of new learnable parameters in the input data for downstream tasks, without altering the pre-trained weights. This method stands out among other lightweight adaptation approaches due to its exceptional adaptability and flexibility [79]. It has garnered increasing attention for adapting vision [26, 27, 80] and vision-language models [28, 33, 81, 82, 30, 83, 84]. Specifically, in VLMs, CoOp [28] pioneers prompt engineering for adapting CLIP models by modeling learnable context vectors to replace hand-crafted text prompts. CoCoOp [33] further enhances the generalization ability of CoOp by introducing conditional prompts specific to each visual input instance. MaPLe [30] integrates vision and language prompts with inner synergy for cross-modality prompt learning. Two recent works, ProGrad [84] and PromptSRC [83], concurrently advance the generalization of prompt learning by employing regulating constraints from zero-shot CLIP predictions to prevent the forgetting of general knowledge.
@@ Appendix C Additional Implementation Details
All experiments are conducted in an environment running PyTorch 1.10.1 and CUDA 11.3 on Python 3.8. Experiments of adversarial prompt tuning on the ImageNet-1K dataset are carried out on a single NVIDIA RTX A40 GPU, while experiments on the other 10 datasets are performed on a single NVIDIA RTX 4090 GPU.
@@ C.1 Additional Implementation Details for Baselines
Adversarial visual prompt. We implement the adversarial visual prompt following all architectural and parameter settings in [11] for a fair comparison. In detail, we follow their code implementation to use a token-level prompt with size 5 and an image padding prompt for 30 pixels around the image. An SGD optimizer and a consine learning rate scheduler are used to train 10 epochs with an initial learning rate of 40.
Adversarial text prompt. We adopt a CoOp architecture [28] as our text prompt baseline and adapt learnable context vectors with adversarial examples. We typically follow [28] to use 16 context tokens with an additional class token appended at the end of the context vector. An SGD optimizer and a consine learning rate scheduler are used to train 200 epochs with an initial learning rate of 0.002, which aligns with the training settings in CoOp.
Adversarial multi-modal prompt. Adversarial multi-modal prompt in this work follows all the design choices as MaPLe [30], but are adapted with an adversarial text-image contrastive loss. To sum up, it contains a token-level learnable token with size 2 in both text and visual branches in the first 9 transformer layers, and the deep prompts are coupled through a text-to-image projection. The above prompt tokens as well as the deep projections are optimized for 10 epochs with SGD optimizer and cosine learning rate scheduler from an initial learning rate of 0.0035.
@@
Overall methodological explanations. We summarize the prompt design and loss function of both baselines and our methods in Table 6. Note that the prompt design for baselines follows the original settings in their corresponding paper, while we replace their loss function with the TeCoA loss for adversarial training and evaluation. This is consistent with the methods used in Mao et al. [11].
@@ Appendix D Additional Experimental Results
D.1 Detailed Results on Adversarial Cross-Dataset Evaluation
For the cross-dataset evaluation, models are adapted on the ImageNet-1K dataset using 16 shots and then assessed for their zero-shot adversarial robustness across 10 distinct datasets, without further downstream tuning. As shown in Table 8, our method outperforms its counterparts in 8/11 datasets and baseline in all 11 datasets. Moreover, it reveals that robust adaptation takes the cost of natural accuracy, as models obtained using various robust adaptation methods exhibit a decline in zero-shot natural accuracy performance on downstream datasets, compared to the original CLIP model.
| Natural Acc (%) | ImageNet-1K | Caltech101 | DTD | EuroSAT | OxfordPets | FGVCAircraft | Food101 | Flowers102 | StanfordCars | SUN397 | UCF101 | Average |
| PGD-100 Acc (%) | ImageNet-1K | Caltech101 | DTD | EuroSAT | OxfordPets | FGVCAircraft | Food101 | Flowers102 | StanfordCars | SUN397 | UCF101 | Average |
D.2 Natural Generalization Gap Hinders Robust Adapting
We identify a failure risk in few-shot adversarial prompt learning using TeCoA loss [11], where insufficient natural generalization on the sampled dataset impedes robustness learning. Figure 6(a) displays the loss variation during training under this setup. Under the same experimental setup using the TeCoA loss, different trials exhibit completely different trends: the curve for the failure case shows that the loss quickly ceases to decline and becomes stable shortly after training begins, whereas the loss in the normal case continues to decrease as the training progresses.
We presume that this failure stems from a lack of natural generalization ability. To confirm this, we first conduct natural tuning on the problematic few-shot dataset and then apply adversarial prompt learning. This restores the model’s robust fine-tuning performance, as evident in Figure 6(b), where natural and robust accuracies improve significantly after natural example adaptation. Besides, we validate the learning process on the same few-shot dataset with a dual-form loss in the training objective that considers both natural and adversarial terms (red lines in Figure 6(a)). It is revealed that this two-term loss effectively acts as a surrogate for the aforementioned two-stage method, avoiding potential failures caused by the natural generalization barrier in end-to-end training.
D.3 Incremental Changes from AdvMaPLe
By examining the structural vulnerabilities (Figure 4) and the inadequate natural generalization (Figure 6) inherent in AdvMaPLe’s learning objectives, we have proposed straightforward yet effective improvements. We generally have two improvements from AdvMaPLe:
Imp.1: Regarding the prompt design, we optimize the projection direction under the adversarial prompt learning situation for superior stability.
Imp.2: Regarding the learning objective, we not only consider the natural generalization gap that may cause the adversarial prompt to fail in the few-shot setting, but also make full use of the CLIP structure to design a differentiated robust learning strategy between different modalities.
For a clear picture of the empirical boost, we demonstrate the incremental changes concerning AdvMaPLe. We separately report the performance changes resulting from modifications to prompt direction alone, the learning objective alone, and the combination of both in Table 9.
We find from Table 9 that adopting our provided learning objective alone can enhance model performance. However, the performance change brought about by modifying the projection direction alone on the basis of AdvMaPLe is subtle, as the model’s performance on most downstream datasets is not saturated at this point. On the other hand, further modifications to the projection direction based on row 3 can lead to additional improvements in model performance due to the repair of instabilities on certain datasets.
D.4 Case Study: Improving AdvVLP with Our Learning Objective